OpenStudy (halfdeafdarling):

Anyone have a good idea on extraneous and not extraneous equations are? I am working on a project and would like help on figuring out what I am needing to do?

OpenStudy (halfdeafdarling):

Hi, looking for help, can you help?

OpenStudy (anonymous):

an extraneous solution is one that you get algebraically, but it is not actually a solution to the equation. What are you trying to do?

OpenStudy (halfdeafdarling):

well I have a three part piece I need to be completing. the first part is this: Create an equation that results in at least one extraneous solution. Work through your equation, justify each step, and explain how the solution is extraneous" I am not sure how to begin with this.

OpenStudy (halfdeafdarling):

i figured out thee third piece though well enough so more so two pieces i need help with --More so i am not sure how to create my own equation

OpenStudy (anonymous):

ok. you sometimes get an extraneous solution when solving a radical equation. Something like this, where c and k are constants \[\sqrt{x \pm c}=x \pm k\]

OpenStudy (anonymous):

for example \[\sqrt{x+2}=x+4\]

OpenStudy (anonymous):

do you know how to solve that?

OpenStudy (anonymous):

ah, that's a nasty equation. Doesn't actually have any real solutions. Make it \[\sqrt{x+2}=x-4\]

OpenStudy (anonymous):

still here?

OpenStudy (halfdeafdarling):

I am I am also going over a khan academy video while reading up on it. First time viewing it in this situation. so let me write down your equation, where would I begin on solving it? Would i square both sides of the equation?

OpenStudy (anonymous):

yes

OpenStudy (halfdeafdarling):

okay I understand that part of squaring so then we would get. \[x+4=x^2-16?\]

OpenStudy (halfdeafdarling):

there is a part where I believe there should be another number at the end, but that is where i get confused. Such as the equation on khan academy I was looking at. \[\sqrt{x}=2x-6\]

OpenStudy (halfdeafdarling):

I get through it to where it is x=4x^2-24x+36. I dont understand on how it is 24 and 36. once they have 12 do they just multiply that why two then multiply the 6 by itself? If so why do they do that?

OpenStudy (anonymous):

You made a mistake squaring the sides. \[(\sqrt{x+2})^2=(x-4)^2\] \[x+2=(x-4)(x-4)\] \[x+2=x^2-8x+16\] Then you get everything to one side

OpenStudy (anonymous):

in the KA example, \[(2x-6)^2 = (2x-6)(2x-6)\] foil to get \[4x^2-24x+36\]

OpenStudy (halfdeafdarling):

I think my mind is making this more difficult than it should be. I see how to get the x^2 but the -8x is still confusing (im understanding the 16 though because the distributing 4*4=16 . the -8x where is that? (sorry if i seem like it is going over my head first time seeing this being done really confused)

OpenStudy (halfdeafdarling):

OOOH is it adding like terms?

OpenStudy (halfdeafdarling):

cause you are getting -4x on each side which would be getting the -8x ?

OpenStudy (anonymous):

It's FOIL/distribution \((x-4)(x-4)=x^2-4x-4x+16=x^2-8x+16\)

OpenStudy (anonymous):

exactly

OpenStudy (halfdeafdarling):

oh okay wow sorry about that, that went over my head there.

OpenStudy (halfdeafdarling):

so then we would have \[x+2=x^2-8x+16\]

OpenStudy (halfdeafdarling):

then we would be subtracting two ?

OpenStudy (anonymous):

yes. subtract 2 and x. You need one side to be 0 to solve the quadratic equation

OpenStudy (halfdeafdarling):

okay got that so then it would be 0=x^2+6

OpenStudy (anonymous):

no. not sure what you did. You actually need to combine like terms.